Tina has 2 credit cards that have a total balance of $32,400. One card has an interest rate of 10% and the other has an interest rate of 8%. The first year accumulated $1892 in interest charges. How much was the loan on each credit card?
Answer by checkley77(12844)

.10x+.08(32,400-x)=1,892
.10x+2,592-.08x=1,892
.02x=1,892-2,592
.02x=-700
x=700/-.02
x=-35,000 does not fit the problem
Test
1,892/32,400=.058395 or 5.84% would be the average percent of the 2 cards.
One month of 8% for the full amount would be:
.08*32,400=$2,592 well above the charges at the end of the first year.
You must have some incorrect figures or facts.